Make New Entry,
Make Followup Entry
User name dalton
Log entry time 15:21:01 on November 22, 2009
Entry number 301845
keyword=Run coordinator daily summary
Run Coordinator's Report for November 21 Day to November 22 Owl
- No beam was taken on swing shift as accelerator first fixed a
cryomodule failure and then a vacuum valve in the South Linac.
- We started slug 7 this morning. Part of the Slug 7 running includes
4 hours of dummy aluminum data.
- A cathode spot move is planned by the injector group either on
Monday or Tuesday.
- A bad scalar was changed and this has corrected the double peaked
asymmetries seen in 3 channels of the right arm pion narrow asymmetries.
- The compton photon detector PMT high voltage is still being run only
up to 1500 V, low energy background cause it to draw too much current.
Some shielding changes were made during the opportunistic access on swing
shift. Studies continue.
- Take production data, including 4 hours of aluminum dummy
data (page Compton experts to do simultaneous low current tests.)
- Monday at 12 pm, begin ramping right dipole to zero current and
field. This is a good time for target boiling studies and compton
- Monday at 3 pm, controlled access - techs will change polarity of
spectrometer magnets. Take reverse polarity data overnight. A minimum
of 5 hours of data is needed here.
- Tuesday at 4 am (if we have 5 hurs of data), begin ramping right
dipole to zero current and field.
- Tuesday at 7, controlled access, techs restore spectrometer
polarity. Resumer normal production running.
- Three shifts need a shift leader: Shift Leader on the day
shift of Wednesday, November 25, 2009 and Shift Leader for the Owl and
Day shifts of Friday, November 27, 2009. Please sign up!
- Shift Leaders: please write a detailed shift summary (including
runs taken and any
activities) and submit the BTA for each shift. Please log CODA reboots
in the halog and shift summary.
- 3rd person: Please make sure to start and stop the HRS DAQ
at the same time as the parity DAQ and to do the prompt analysis and the
HRS analyzer for every run.
- 3rd person: please keep a careful eye on the online data display.
Most problems can be caught quickly this way.
- Please make sure that the shift checklist is done once each shift.
- Production charge taken in the 24 hours ending Sunday Owl: 2.3 C
- Total production charge taken in all, end of Sunday Owl: 24.0 C
Shift Summaries and BTA Reports